1. At what temperature will 0.654 moles of neon gas occupy 12.30 L at 1.95 atm?
  2. A sample of argon gas at STP occupies 56.2 L. Determine the number of moles of argon.
  3. A gas is compressed in a 25 L storage tank under a pressure of 150 kPa and a temp of
  4. 21°C. How many moles of gas are contained in the tank?
  5. What volume will 45 grams of hydrogen gas occupy at 1.05 atm and 25°C?
  6. At what temperature will 5.00 g of Cl2 exert a pressure of 900 mm Hg at a volume of 750.0 mL?
  7. Ammonium nitrate decomposes explosively upon heating according to the following equation:
  8.             NH4NO3(s)  N2  (g) + O2 (g) + H2O (g) 
  9. Calculate the total volume of gas produced at 125°C and 748 mm Hg when 1.55 kg of ammonium nitrate completely decomposes. Balance the equation! 




1
Expert's answer
2021-02-17T07:47:46-0500

Q164327


1) At what temperature will 0.654 moles of neon gas occupy 12.30 L at 1.95 atm?

Solution:

T = unknown, n = 0.654 mol, V = 12.30 L, P = 1.95 atm .


We can use the ideal gas equation, PV = nRT, and find the temperature of the gas.


"PV=nRT ;" where R = 0.08206 L-atm/mol.K


substitute all the given information in the ideal gas equation we have


"1.95 atm.12.30 L = 0.654mol * 0.08206L.atm\/mol.K*T"


"\frac{23.985 L.atm }{0.05367L.atm/K}= \frac{0.05367L.atm/K*T }{0.05367L.atm/K}"


divide both the side by 0.05367 L.atm/K we have


T = 446.9K. = 446.9K - 273.15 = 173.8 °C


Hence the neon gas will be at 447 Kelvin temperature.

which in degree celsius will be 174 °C



2) A sample of argon gas at STP occupies 56.2 L. Determine the number of moles of argon.

Solution:

At STP, T = 273.15K and P = 1atm.

We are given V = 56.2 L.

Gas constant, R = 0.08206 L.atm/mol.K.


Substitute all this information in the ideal gas equation, PV = nRT we have



"PV = nRT"


"1atm * 56.2L = n * 0.08206 L.atm\/mol.K* 273.15K."


"56.2L.atm = n * 22.415 L.atm\/mol"


divide both the side by 22.415L.atm/mol, we have


"\frac{56.2L.atm}{ 22.415 L.atm/mol }=\frac{ n * 22.415 L.atm/mol }{22.415 L.atm/mol}"


n = 2.5073 moles of Argon.


in the question, we are given 56.2 L in 3 significant figure, so our final answer must also be in

3 significant figure.


Hence there would be 2.51 moles of Argon present in the given sample.



3) A gas is compressed in a 25 L storage tank under a pressure of 150 kPa and a temp of

21°C. How many moles of gas are contained in the tank?


Solution:

We are given Volume, V = 25L, Pressure, P = 150 kPa

Temperature T = 21°C.


In the ideal gas equation, pressure must be used in the 'átm' unit and temperature must be in the Kelvin unit.


"Pressure \space P \space in\space atm = 150 kPa * \frac{1atm }{101.325 kPa } = 1.4804 atm."


"Temperature\space in\space Kelvin = 21°C = 21 + 273.15 = 294.15K"


Gas constant, R = 0.08206 L.atm/mol.K

substitute all this information in the ideal gas equation we have



"PV = nRT"



"1.480atm * 25L = n * 0.08206 L.atm\/mol.K * 294.15K"


"37L.atm = n * 22.415 L.atm\/mol"



divide both the side by 22.415L.atm/mol, we have



"\frac{37L.atm}{ 22.415 L.atm/mol }=\frac{ n * 22.415 L.atm/mol }{22.415 L.atm/mol}"



"n = \frac{37L.atm}{ 22.415 L.atm/mol }"


n = 1.6507 moles.


in the question, we are given pressure and volume both in 2 significant figure, so our final answer must

also, be in 2 significant figures.


Hence there would be 1.7 moles of gas in the tank.




Question # 4 is not present



5) What volume will 45 grams of hydrogen gas occupy at 1.05 atm and 25°C?


Solution:


Here instead of 'moles of hydrogen gas' we are given the grams of hydrogen gas.

So we will first convert 45 grams of H2 gas to moles by using the molar mass of H2 gas.


molar mass of H2 = 2 * atomic mass of H = 2 * 1.00794 g/mol

= 2.016 g/mol



moles of H2 "= 45 \space g \space of \space H_2 * \frac{1 \space mol \space H_2}{2.016 \space g of \space H_2}" "= 22.32\space mol\space of \space H_2"


So now we have

n = 22.32 moles , P = 1.05 atm .

T = 25°C = 25 + 273.15 = 298.15K.

Gas constant, R = 0.08206 L.atm/mol.K


substitute this information in the ideal gas equation we have



"PV = nRT"


1.05 atm * V = 22.32 mol * 0.08206 L.atm/mol.K * 298.15K


1.05 atm * V = 546.09 L.atm.


divide both the side by 1.05 atm, we have


"\frac{1.05 atm * V }{1.05 atm }= \frac{546.09 L.atm. }{1.05 atm }"


V = 520.08 L.


In the given question the quantity with the least number of the significant figure is 45 grams.

It is in 2 significant figures. So our final answer must also be in 2 significant figures.


Hence the volume occupied by the hydrogen gas will be 520 L.




6) At what temperature will 5.00 g of Cl2 exert a pressure of 900 mm Hg at a volume of 750.0 mL?

Solution.

Here too we are given a mass of Cl2 gas. So first we will have to convert it to grams by using

the molar mass of Cl2.


molar mass of Cl2 = 2 * atomic mass of Cl = 2 * 35.453 g/mol

= 70.906 g/mol



moles of Cl2 "= 5.00 g of Cl_2 * \frac{1 mol\space of\space Cl_2 }{70.906\space g\space\space of\space Cl_2 }" "= 0.07052 \space moles\space of\space Cl_2"



Convert 750.0mL to 'L' and 900 mm Hg to 'atm' .


"volume\space in\space 'L' = 750.0mL * \frac{1L}{1000mL} = 0.750 L"


"pressure\space in\space 'atm' = 900 mm Hg * \frac{1atm }{760 mm Hg } = 1.184 atm." ( since, 1atm = 760 mm Hg)



So now we have

moles of Cl2 = 0.07052 mol

volume occupied by Cl2 = 0.750 L

pressure = 1.184 atm.

gas constant, R = 0.08206 L.atm/mol.K


plug all this information in the idea gas equation we have



"PV = nRT"




"1.184 atm * 0.750L = 0.07052 mol * 0.08206 L.atm\/mol.K * T."


0.8882 L.atm = 0.005787 L-atm/K * T


divide both the side by 0.005784 L.atm/mol, we have


"\frac{0.8882 L.atm}{0.005787 L-atm/K} = \frac{0.005787 L-atm/K * T }{0.005787 L-atm/K}"



"T =\frac{0.8882 L.atm}{0.005787 L-atm/K} = 153.48 K ."


which in 3 significant figure is 153 K or in

degree celsius is = 153.45 K - 273.15 = -119.66 °C = -120 °C



Hence the temperature of the Cl2 gas will be 153 K or -120 °C.




7) Ammonium nitrate decomposes explosively upon heating according to the following equation:

            NH4NO3(s) ----->  N2  (g) + O2 (g) + H2O (g) 

Calculate the total volume of gas produced at 125°C and 748 mm Hg when 1.55 kg of ammonium nitrate completely decomposes. Balance the equation! 


Solution:


Our first step will be to balance the given reaction.


 NH4NO3(s) ----->  N2  (g) + O2 (g) + H2O (g) 


We will first balance the element which is not present in a single elemental form.

So N and O will be balanced afterward. First, we will balance hydrogen.


There is 4 H on the left side and 2 on the right side, so to balance the H write 2 in front of H2O.

This will balance H on both the side.


NH4NO3(s) ----->  N2  (g) + O2 (g) + 2 H2O (g) 


Nitrogen is already balanced here.

There are 3 O on the left side and 4 on the right side.


to balance the oxygen we will have to write a proper coefficient in front of O2.

placing a fraction 1/2 in front of O2 will make oxygen equal on both the side.


NH4NO3(s) -----> N2  (g) + 1/2 O2 (g) + 2 H2O (g) 


We can't keep fractions in our balanced equation. To remove fraction multiply the whole reaction by 2.


2 * ( NH4NO3(s) -----> N2  (g) + 1/2 O2 (g) + 2 H2O (g) )


2 NH4NO3(s) ----->  2 N2  (g) + 1 O2 (g) + 2 H2O (g)


So now this reaction is balanced.



Step 1: Convert 1.55 kg of NH4 NO3 to moles by using the molar mass of NH4NO3



1.55 kg NH4NO3 in grams = 1.55 kg * 1000g/1kg = 1550 grams of NH4NO3



molar mass of NH4 NO3 = 2 * atomic mass of N + 3 * atomic mass of O + 4 * atomic mass of H

= 2 * 14.007 g/mol + 3 * 15.999 g/mol + 4 * 1.00794 g/mol

= 80.0426g/mol



moles of NH4NO3 = 1550 g of NH4NO3 * 1mol NH4NO3 / 80.0426 g NH4NO3

= 19.365 mol NH4NO3




in the balanced reaction, we can see that


2 NH4NO3(s) ----->  2 N2  (g) + 1 O2 (g) + 2 H2O (g)


2 moles of NH4 NO3 gives total = 2 + 1 + 2 = 5 moles of gas.


so 19.365 moles of NH4NO3 will give


= 19.365 mol NH4NO3 * 5 mol gas / 2 mol NH4 NO3 = 48.412 mol gas



Step 2: To find the volume occupied by the gas.


We have moles of gas = 48.412 moles


T = 125°C = 398.15K

Pressure P = 748 mm Hg, which in átm = 748mmHg * 1atm / 760 mm Hg = 0.984 atm .


Volume = unknown

gas constant R = 0.08206 L.atm/mol.K

moles of gas = 48.412 mol


plug all this information in the ideal gas equation, PV = nRT


0.984 atm * V = 48.412 mol * 0.08206 L.atm/mol.K * 398.15K


0.984atm * V = 1581.7 L-atm


divide both the side by 0.984 atm we have


V = 1581.73L-atm/ 0.984 atm = 1607.4 L



Hence the volume occupied by the combusted gas after completely decompostion of NH4NO3 is

1600Liters.